If Ø(x, y, z) = x² - y²-2²-2, then ∇Ø at (2,1,-1) is


A) 4i +2j-2k

B)4i-2j-2k

C) 4i + 2j + 2k

D)4i - 2j +2k


1
Expert's answer
2021-08-16T02:19:54-0400

"Ø(x, y, z) = x² - y²-z²-2\\
Then \space ∇Ø= \frac{\partial}{\partial x}i, \frac{\partial}{\partial x}j, \frac{\partial}{\partial x}k\\
∇Ø= \frac{\partial}{\partial x} (x² - y²-z²-2), \frac{\partial}{\partial x}(x² - y²-z²-2), \frac{\partial}{\partial x}(x² - y²-z²-2)\\
∇Ø=2x i, 2yj,2zk\\
 At (2,1,-1) \\
∇Ø=4 i+ 2j-2k"


Option A is correct.